OpenStudy (anonymous):

Solve the equation for a. K = 4a + 9ab 1. a = K(4 + 9b) 2. a = 4(K + 9b) 3. a = 4+9b/K 4. a = K/ 4+9b

OpenStudy (anonymous):

a=0

OpenStudy (anonymous):

choices are : 1. a = K(4 + 9b) 2. a = 4(K + 9b) 3. a = 4+9b/K 4. a = K/ 4+9b

OpenStudy (paki):

@nandabo_ step 1: take "a" common.... K = 4a + 9ab k = a(4+9b) step 2: leave "a" alone now.... so by doing so just divide this equation with (4+9b) a = k/(4+9b)...... and that is your required....